# Extracting the Speed of Sound in Heavy-Ion Collisions: A Study of Quantum Fluctuations and Thermalization

### v1

@@ Abstract
Importantly, for a thermalized QGP, $\delta$ is independent of the thermodynamic response relation. This independence reflects the quantum origin of $\delta$, which introduces stochastic fluctuations uncorrelated with the deterministic macroscopic relationship. While $\delta$ is independent of the thermodynamic response, it may exhibit statistical correlations with individual observables like $\Delta_{N}$ or $\Delta_{p}$, arising from shared quantum or initial-state fluctuations. In a thermalized system, where $\delta$ varies randomly and independently across events, its distribution approaches Gaussianity by the Central Limit Theorem (CLT) Samanta et al. (2024, 2023). On the other hand, in a system without thermalization, the response relation itself (where the response coefficient is no longer the speed of sound) is quantum, hence $\delta$ is not independent from event to event and the $\delta$-distribution is generically non-Gaussian.

### v2

@@ Abstract
Importantly, for a thermalized QGP, $\delta$ is independent of the thermodynamic response relation. This independence reflects the quantum origin of $\delta$, which introduces stochastic fluctuations uncorrelated with the deterministic macroscopic relationship. While $\delta$ is independent of the thermodynamic response, it may exhibit statistical correlations with individual observables like $\Delta_{N}$ or $\Delta_{p}$, arising from shared quantum or initial-state fluctuations. In a thermalized system, where $\delta$ varies randomly and independently across events, its distribution approaches Gaussianity by the Central Limit Theorem (CLT) Samanta et al. (2024, 2023). On the other hand, in a system without thermalization, the response relation itself (where the response coefficient is no longer the speed of sound) is quantum, hence $\delta$ is not independent from event to event and the $\delta$-distribution is generically non-Gaussian. This non-Gaussinity in $\delta$ in non-thermal systems can be conceptually understood as a consenquence of interactions among these fluctuations, through the quantum response relation. This is analogous to the non-Gaussianity observed in primordial quantum fluctuations, which signals interactions in the early universe Maldacena (2003).
@@
Supplemental Material
The fluid-like fireball created in heavy-ion collisions converts to particles on a hyper-surface $\Sigma$, subject to the freeze-out condition. Through this conversion, the energy and entropy of the expanding fireball are expressed as
[…]
where $\sigma^{\mu}$ denotes the normal vector to the freeze-out hyper-surface, and $T^{\mu\nu}$ and $s^{\mu}$ are the relativistic energy-momentum tensor and entropy flow vector of the fluid, respectively. With respect to the locally equilibrated equation of state (LEOS) that relates energy or entropy density to temperature, the fireball can be effectively reduced to a uniform system characterized by the effective temperature $T_{\rm eff}$ and effective volume $V_{\rm eff}$, as described by the second equation in Eq. (6).
Particles generated from the freeze-out hyper-surface evolve further according to the UrQMD model. The resulting mean transverse momentum $\langle p_{T}\rangle$ of these final particles is expected to exhibit a linear correlation with the effective temperature $T_{\rm eff}$ on an event-by-event basis. For Pb-Pb collisions, which are approximately 2+1 dimensional due to longitudinal boost invariance, we find that $\langle p_{T}\rangle^{(i)}\approx 3T_{\rm eff}^{(i)}$, where $i$ labels the event.
In the p-Pb system, however, longitudinal boost invariance no longer holds, introducing ambiguity in the total energy determined by Eq. (6). The effective temperatures obtained using the $(t,x,y,z)$ coordinate system and the $(\tau,x,y,\eta_{s})$ Milne coordinate system—where $\tau=\sqrt{t^{2}-z^{2}}$ and $\eta_{s}=\tanh^{-1}(z/t)$—are found to differ. Despite this discrepancy, solutions from both coordinate systems exhibit a strong linear correlation with $p_{T}^{(i)}$ on an event-by-event basis, although the correlation coefficients differ. This difference implies that, in a fireball without boost invariance, the mean transverse momentum of the final-state particles may receive contributions from the kinematic energy associated with longitudinal expansion. As there are two sets of linear relations to the effective temperatures, we consider the realistic energy scale of the fireball that is characterized by the mean transverse momentum in the extraction of the speed of sound lies in between, which gives rise to a range in the linear relation between $\langle p_{T}\rangle^{(i)}$ and $T_{\rm eff}^{(i)}$.